# Text-Phase Synergy Network with Dual Priors for Unsupervised Cross-Domain Image Retrieval

### 13 Mar 2026 v1

@@ 2 Related Work
Cross-domain image retrieval involves retrieving images where the query and gallery sets originate from different domains. Existing methods [15, 16, 24] predominantly rely on supervised learning with labeled data, which limits scalability in real-world applications. To address this limitation, recent research has focused on the setting of unsupervised cross-domain image retrieval (UCDIR), which seeks to learn domain-invariant semantic representations from unlabeled data. Representative approaches such as DD [13] and CODA [42] employ $K$-means clustering to generate pseudo-labels. DD further utilizes these pseudo-labels for intra-domain contrastive learning and minimizes the Distance-of-Distance loss to align distributions, while CODA leverages pseudo-labels to initialize both intra-domain and cross-domain classifiers, enabling the separate learning of intra-domain representations and cross-domain feature alignment. However, these methods decompose the UCDIR task into two independent stages, intra-domain representation learning and cross-domain alignment, overlooking the intrinsic correlation between them. To address this, ProtoOT [20] and ShieldIR [34] unify the two stages within an optimal transport framework, combining prototype alignment with contrastive learning. However, semantic degradation persists due to the entanglement of domain-specific and semantic features. SA-MoE [38] addresses this issue by emphasizing semantic modeling through semantically attentive fusion and contextual relevance, thereby preserving semantic integrity and underscoring the importance of semantics in UCDIR.
@@ 3.2 Domain Prompt Generation Module
where $I_{i}$ and $T_{i}$ denote the re-paired image and text embeddings based on cosine similarity scores, which allows the model to partially correct inaccurate pseudo-labels through contrastive alignment. $s(I_{i},T_{i})=\frac{I_{i}^{\top}T_{i}}{\|I_{i}\|\|T_{i}\|}$ and $\tau$ is the temperature. $N$ is the batch size, $P(y_{i})$ is the set of positive samples with respect to $T_{y_{i}}$, i.e., $P(y_{i})=p\in\{1...N|y_{p}=y_{i}\}$. The symbol $|\cdot|$ denotes the Cardinality of this set. Then the proposed DPG is optimized via the following loss: $\mathcal{L}_{prompt}=\mathcal{L}_{i2t}^{i}+\mathcal{L}_{t2i}^{i}$.
@@ 4.1 Datasets and Setting
We evaluate the proposed TPSNet on two widely used datasets for cross-domain retrieval: Office-Home and DomainNet. Office-Home dataset [37] comprises 65 classes and 4 domains (Art, Clipart, Product, Real). DomainNet dataset [29] consists of 6 domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch. Following the criteria established by DD [13], we select 7 classes from the DomainNet dataset with over 200 images in each domain for training and testing.
@@
Following DD [13], we assess retrieval performance on the Office-Home dataset using precision at top-$k$ retrieved images, specifically at $k=1$, 55, and $15$ (i.e., $P@1$, $P@5$, and $P@15$). For the DomainNet dataset, we report precision at $k=50$, 100100, and $200$ (i.e., $P@50$, $P@100$, and $P@200$).
@@ 4.2 Comparison with State-of-the-Art Methods
In our experimental validation, we evaluate the effectiveness of TPSNet by comparing it with the following methods: DD [13], CoDA [42], DG-UCDIR [14], ProtoOT [20], ShieldIR [34] and SA-MoE [38]. Table 1 and Table 2 show the detailed retrieval performance for Office-Home and DomainNet, respectively. TPSNet’s impressive performance is evident from the summarized results in Figure 3, with detailed average accuracies provided in Table 3.
@@ Appendix B Additional Experiments
In addition, we compare TPSNet with the current state-of-the-art large multimodal models (LMMs) for retrieval performance, including InternVL3.5 [41], SAIL-VL2 [45], MM-Embed [23], Qwen2.5VL [2], and LamRA [25]. The average retrieval performance is summarized in Table 12. Since InternVL3.5 and SAIL-VL2 are primarily designed for question-answering tasks, we use only their pre-trained vision encoders. All other methods utilize the complete output vectors of LMMs as feature representations. Experimental results demonstrate that, although LMM-based universal retrieval models exhibit strong zero-shot generalization ability, there remains a noticeable performance gap compared to TPSNet, particularly in scenarios involving significant domain discrepancies. Specifically, their performance deteriorates significantly in domains with substantial background noise and abstract visual patterns, such as Infograph and Quickdraw. This degradation can be attributed to the inherent bias of foundation models towards natural-image domains prevalent in their pretraining datasets, which limits their ability to generalize to highly stylized domains.

### 03 Apr 2026 v2

@@ 2 Related Work
Cross-domain image retrieval involves retrieving images where the query and gallery sets originate from different domains. Existing methods [15, 16, 24] predominantly rely on supervised learning with labeled data, which limits scalability in real-world applications. To address this limitation, recent research has focused on the setting of unsupervised cross-domain image retrieval (UCDIR), which seeks to learn domain-invariant semantic representations from unlabeled data. Representative approaches such as $D$ [13] and CODA [42] employ $K$-means clustering to generate pseudo-labels. $D$ further utilizes these pseudo-labels for intra-domain contrastive learning and minimizes the Distance-of-Distance loss to align distributions, while CODA leverages pseudo-labels to initialize both intra-domain and cross-domain classifiers, enabling the separate learning of intra-domain representations and cross-domain feature alignment. However, these methods decompose the UCDIR task into two independent stages, intra-domain representation learning and cross-domain alignment, overlooking the intrinsic correlation between them. To address this, ProtoOT [20] and ShieldIR [34] unify the two stages within an optimal transport framework, combining prototype alignment with contrastive learning. However, semantic degradation persists due to the entanglement of domain-specific and semantic features. SA-MoE [38] addresses this issue by emphasizing semantic modeling through semantically attentive fusion and contextual relevance, thereby preserving semantic integrity and underscoring the importance of semantics in UCDIR.
@@ 3.2 Domain Prompt Generation Module
where $I_{i}$ and $T_{i}$ denote the re-paired image and text embeddings based on cosine similarity scores, which allows the model to partially correct inaccurate pseudo-labels through contrastive alignment. $s(I_{i},T_{i})=\frac{I_{i}^{\top}T_{i}}{\|I_{i}\|\|T_{i}\|}$ and $\tau$ is the temperature. NN is the batch size, $P(y_{i})$ is the set of positive samples with respect to $T_{y_{i}}$, i.e., $P(y_{i})=p\in\{1...N|y_{p}=y_{i}\}$. The symbol $|\cdot|$ denotes the Cardinality of this set. Then the proposed DPG is optimized via the following loss: $\mathcal{L}_{prompt}=\mathcal{L}_{i2t}^{i}+\mathcal{L}_{t2i}^{i}$.
@@ 4.1 Datasets and Setting
We evaluate the proposed TPSNet on two widely used datasets for cross-domain retrieval: Office-Home and DomainNet. Office-Home dataset [37] comprises 65 classes and 4 domains (Art, Clipart, Product, Real). DomainNet dataset [29] consists of 6 domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch. Following the criteria established by $D$ [13], we select 7 classes from the DomainNet dataset with over 200 images in each domain for training and testing.
@@
Following $D$ [13], we assess retrieval performance on the Office-Home dataset using precision at top-$k$ retrieved images, specifically at $k=1$, 55, and 1515 (i.e., $P@1$, $P@5$, and $P@15$). For the DomainNet dataset, we report precision at $k=50$, 100100, and 200200 (i.e., $P@50$, $P@100$, and $P@200$).
@@ 4.2 Comparison with State-of-the-Art Methods
In our experimental validation, we evaluate the effectiveness of TPSNet by comparing it with the following methods: $D$ [13], CoDA [42], DG-UCDIR [14], ProtoOT [20], ShieldIR [34] and SA-MoE [38]. Table 1 and Table 2 show the detailed retrieval performance for Office-Home and DomainNet, respectively. TPSNet’s impressive performance is evident from the summarized results in Figure 3, with detailed average accuracies provided in Table 3.
@@ Appendix B Additional Experiments
In addition, we compare TPSNet with the current state-of-the-art large multimodal models (LMMs) for retrieval performance, including InternVL3.5 [41], SAIL-VL2 [45], $M$-Embed [23], Qwen2.5VL [2], and LamRA [25]. The average retrieval performance is summarized in Table 12. Since InternVL3.5 and SAIL-VL2 are primarily designed for question-answering tasks, we use only their pre-trained vision encoders. All other methods utilize the complete output vectors of LMMs as feature representations. Experimental results demonstrate that, although LMM-based universal retrieval models exhibit strong zero-shot generalization ability, there remains a noticeable performance gap compared to TPSNet, particularly in scenarios involving significant domain discrepancies. Specifically, their performance deteriorates significantly in domains with substantial background noise and abstract visual patterns, such as Infograph and Quickdraw. This degradation can be attributed to the inherent bias of foundation models towards natural-image domains prevalent in their pretraining datasets, which limits their ability to generalize to highly stylized domains.